Project

General

Profile

11122013

Nightly Observing Plan for MJD 56609 (starting evening of Nov 12, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Matt Wiesner Liz Buckley-Geer Marisa March
ext. 515 ext. 512 ext. 516 or 630 338 9267

CTIO staff (Telescope Operators)

Day Support Night Assistant
Hernan Tirado Leonardo Paredes

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56609, date 20131112, Site CTIO, UT-offset 3 hr
Current local time 14:56 CLST     UT 17:56     SidTim 16:42
Sunset     20:22 CLST  23:22 UT   SidTim  332 deg  
Eve. -10   21:00 CLST  00:00 UT   SidTim  342 deg  Zeropoint; standards
Eve. -14   21:22 CLST  00:22 UT   SidTim  347 deg  Begin obstac
Eve. -18   21:44 CLST  00:44 UT   SidTim  352 deg  Last chance standards
Midpoint   01:27 CLST  04:27 UT   SidTim   48 deg  
Moonset    03:59 CLST  06:59 UT   SidTim   86 deg  
Morn -18   05:10 CLST  08:10 UT   SidTim  104 deg  Begin astronomical twilight
Morn Std   05:32 CLST  08:32 UT   SidTim  110 deg  End obstac; begin standards
Morn -10   05:53 CLST  08:53 UT   SidTim  115 deg  
Sunrise    06:32 CLST  09:32 UT   SidTim  125 deg  
Moon position (deg): RA 356.4, Dec 2.9 Phase: 0.703

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Clear

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards
  3. ObsTac
  4. End of night standards
  5. End of night calibrations

Start of Night Standards
UT Date = '2013/11/13 00:00' Local time = 21:00
SDSSJ0200-0000 ra 02:00:00 dec +00:00:00 secz 1.7 script to use: DES_std_sdssj0200_Yzirg.json
SDSSJ0100-0000 ra 01:00:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj0100_Yzirg.json
SDSSJ2140-0000 ra 21:40:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj2140_Yzirg.json

End of Night Standards
UT Date = '2013/11/13 08:32' Local time = 05:32
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.9 script to use: DES_std_sdssj1048_grizY.json
SDSSJ0958-0010 ra 09:58:00 dec -00:10:00 secz 1.5 script to use: DES_std_sdssj0958_grizY.json
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY.json

Notes from the 4PM Meeting and Additional Comments:

Run Manager's Meeting 11/12/13
------------------------------
Marisa, Alex, Eric, Matt, Tom, Liz, Hernan, Manual, Brian, Josh
Robert, Margaret, Klaus

Telescope and Camera
--------------------
Chiller #1 failed last night at 3:15ish local, just before the
Moon was about to set. The camera is operating on chiller #2.
NESLab is running a little cooler because the flow is lower.
Tomorrow the chiller will be repaired by the Trane people.

Telescope is ready to start calibrations. SISPI is restarted and
ready to go.

TASCA was not running last night. Team to check with Hernan.

Last Night's Data
-----------------
Median seeing i-band zenith was 0.9", nice night.
DIMM was reporting 0.55" zenith g-ish. Got a lot of z-band
images because the Moon was up. But we didn't get the highest
priority z-band area because the Moon was too close to it.

Image 253214 is bad because of temperature of focal plane.
Images after that should be ignored.

DESDM
-----
Network problems wrecked processing from 11/9,10.
But they did get all the data.

Marisa is putting SN DQ in DESDM data base field SN-Quality.

DQ report from SK on the SN data from 11/9: All Bad. Eric to implement that
into the exposure table.
So we need the script that takes the quality info from the
DESDM fields and puts it into the exposure table. Robert is doing that.
But we are not ready to use the wide-band data DQ because
of the new extinction calculation. So we are not ready to use this
database.

For Tonight
-----------
Weather tonight looks good.

Update in DQ puts us likely to do SN. Otherwise z-band WF.
Obstac change from
1) SN deadman
2) WF grizY if seeing better than 1.1" and fits in the strict tau cut
3) SN tactitcian for fields > 3 days
4) g or Y-band unless seiing > 10"
5) SNX3

changes to

1) SN deadman
2) WF grizY if seeing better than 1.1" and fits in the strict tau cut
3) SN tactician for fields > 3 days
3.5) if seeing better than 1.6" grzY if it fits in a relaxed tau cut (based on
airmass, sky background, seeing)
4) g, or Y-band unless seeing > 10"
5) SNX3

if Eric can swing it.

AOB
---
Last night there were a couple 4 or 5 image blips of bad seeing that resulted in
setting up SN observing under condition 3). Conditions quickly went back to condition
2) but the latency meant 45 minutes or so was spent under presumption of worse
condition. Observers asked if they should have stopped the queue and clear out the
upcoming observations so that Obstac could redecide to do 2).

No instructions exist for this.
Here we try to put down some principals, at least.
0) That Obstac has no internal procedure to "change its mind" is a feature
a) Run Managers decide this
b) It should not be done very often.
c) If done, it should be done so that it doesn't waste partially completed
Supernovae fields.
d) If done right and the gamble pays off ...